ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 28 Vlad, Carmen Doina, Pace University, New York, USA Outer measures and countably additivity of measures Consider X an arbitrary set, L a lattice of subsets of X and A(L) the algebra generated by L. The set M(L) consists of those non negative, finite, finitely additive measures on A(L). For a measure m of M(L), conditions in terms of an induced measure on a Wallman space are given for m to be countably additive on L or on A(L). Relationships between regularity, countably additivity and measurability are investigated. Reference: C. Vlad, Remarks on finitely subadditive outer measures, Annales des Sciences Mathematiques du Quebec, 20 ( 1996 ) no. 1, 93-103.